# Time as a Cosmological Phenomenon

### v2

@@ 1 Introduction
  • Geometrical reasons: The universe does not admit a well defined notion of cosmic time, or, if it does, it started in a high-entropy state, i.e. either Weyl’s postulate [24] or the past hypothesis [25] (or both) are violated. Geometrical obstructions of this type prevent even a local arrow of time from being defined.
  • Topological reasons: The universe is not time-orientable [26] or admits closed timelike curves (CTC) [27] that violate causality and along which entropy cannot globally increase, i.e. the universe is not chronological [28]. Topological obstructions of this type prevent a global arrow of time from being defined everywhere on the spacetime manifold.
@@ 3 The Past Hypothesis
What further constraint does this place on the metric (2.3)? As we saw, the metric admits a preferred slicing with timelike four-velocity field $u^{\mu}=(1,0,0,0)$, therefore the SEC gives
$R_{\mu\nu}u^{\mu}u^{\nu}=R_{00}\geq 0.$ (3.7)
The condition $R_{00}\geq 0$ effectively selects a subset of all possible cosmic-time metrics (2.3). To see this, let us consider the simple case of a homogeneous and isotropic universe. The line-element is given then by the FRW metric:
@@
where $a(t)>0$ is the scale factor and k=1,0,1k=-1,0,1 the curvature parameter. Now, one can show that for the metric (3.8), $R_{00}=-3\ddot{a}/a$, therefore $R_{00}\geq 0$ whenever $\ddot{a}\leq 0$ for all values of $t$ (the universe is always decelerating). In fact, for a perfect fluid source that obeys the SEC, this is always the case and $a(t)\propto t^{n}$ with $0<n<1$.
Exotic matter or quantum effects that violate the SEC can give rise to cosmological models that have no beginning. For example, the simple bouncing model described by
$a(t)=a_{b}\left[1+\left(\frac{t}{t_{b}}\right)^{2}\right]^{q},\quad q>0,$ (3.9)
can be shown to arise due to quantum cosmological effects described by the Wheeler-DeWitt equation [41] and has no cosmological singularity. In fact, in this model $\ddot{a}\leq 0$ only when $|t/t_{b}|\geq 1/\sqrt{1-2q}$ with $0<q<1/2$. This means that around the bounce, i.e. for $|t/t_{b}|\leq 1/\sqrt{1-2q}$, the universe reverses course, starts expanding, and the singularity is avoided [^16] A similar period of rapid expansion happens in inflationary cosmology, which indeed also violates the SEC. However, inflationary spacetimes still exhibit a cosmological singularity (or at least a spacetime boundary): the Borde-Guth-Vilenkin (BGV) theorem shows that any spacetime whose average Hubble expansion rate is positive is geodesically incomplete even when the SEC is violated [42]. The average Hubble rate in the bouncing model of (3.9) on the other hand integrates to zero, so the BGV condition, and the singularity, may be avoided..
@@
$g_{ij}(t,\vec{x})=a^{2}(t)\left[\gamma_{ij}^{(0)}(\vec{x})+t^{n}\gamma_{ij}^{(1)}(\vec{x})+...\right],\quad n>0,$ (3.11)
@@
To summarize, if $\omega_{\mu\nu}=0$ and $R_{00}\geq 0$ everywhere on the manifold $\mathcal{M}$, and $g_{ij}(t,\vec{x})\sim a^{2}(t)\gamma_{ij}^{(0)}(\vec{x})$ for $t\rightarrow 0$, the universe admits a notion of cosmic time with a well-defined direction (that orthogonal to the hypersurfaces of constant cosmic time) and orientation (proceeding from the low-entropy initial singularity). While this ensures the existence of an arrow of time on a local patch, there could still be topological effects that spoil chronology globally, as we will see in the next two sections.
@@ 4 Time-orientability
$(t,\vec{x})\sim\left((-1)^{n}t,\vec{x}+n\vec{a}\right)\quad\forall n\in\mathbb{Z}.$ (4.3)
@@ 5 Chronology
$\gamma:\quad r=r_{0}=\text{const},\quad z=\text{const},\quad\tau=\tau_{0}=\text{const},\quad\phi\in[0,2\pi].$ (5.5)
@@ 6 The Role of Entropy
The first describes an expanding universe, while the second describes a contracting universe. Both are perfectly legitimate solutions of the dynamical equations, but only one of them can be physically realized. Whichever one chooses, time-reversal invariance will be broken.
@@
In §3 we argued that the entropy requirement can be reformulated as a geometric condition: the Weyl curvature hypothesis. This does not contradict the idea that the arrow of time arises from the physical law of entropy increasing in a single direction. Rather, it reflects the fact that the entropy of the gravitational field is inherently geometric, since gravity itself is described by a geometric theory. From this perspective, time should be regarded as a cosmological phenomenon: it is fundamentally tied to the large-scale geometry and topology of the universe, which must allow for a global asymmetry in the cosmic time direction. However, it is not just any geometric asymmetry of spacetime that will yield an arrow of time, but only the particular asymmetry encoded in one specific geometric object (the Weyl curvature tensor) which has a direct link to the entropy of the gravitational field.
There is another reason to believe that entropy must enter somehow into the definition of the arrow of time and this has to do with memory [11, 56]: the formation, stability, and accessibility of memories, whether in physical records, biological systems, or cognitive processes, require irreversible dynamics driven by an entropy gradient. The very possibility of encoding and retrieving information about past states depends on the dissipation of free energy into heat, so that in a perfectly reversible world, no memory could persist. This ties the thermodynamic arrow directly to the psychological arrow of time: our perception of temporal flow, grounded in the accumulation of records of the past but not the future, emerges from the same entropic asymmetry that shapes the universe at large. Without increasing entropy, neither physical traces nor subjective experience of time’s direction could exist [^22] Castagnino and Lombardi argue instead that, once a global temporal asymmetry is found, the past-future distinction is conventional and can be expressed locally as a four-dimensional energy flow, without invoking entropy. We agree this is a coherent view, but maintain that entropy is essential to connect global asymmetry with irreversibility, memory, and the experienced passage of time..
@@ 7 The Problem of Time
$\mathcal{C}=0,\quad\mathcal{C}^{i}=0.$ (7.13)
@@
$K_{ij}=\frac{1}{2N}\dot{h}_{ij}=\frac{a\dot{a}}{N}\gamma_{ij},\quad K=\frac{3\dot{a}}{Na}.$ (7.19)
@@
$\pi^{ij}=\frac{-\sqrt{h}\dot{a}}{N\kappa a}h^{ij},\quad\pi=-\frac{3\sqrt{h}\dot{a}}{N\kappa a}.$ (7.20)
@@ 8 Conclusions
Picking a preferred direction of time is only possible under very special circumstances, namely whenever the universe obeys certain geometrical and topological conditions:
@@
According this view, what we understand as time (change, flux, becoming, etc.) is not tied in any way to fundamental physics, but is instead a product of the distribution of matter and energy in the whole universe, or, equivalently, since matter and energy determines geometry, of its shape. A simple example is given by irreversibility, which is one of the basic features of our experience of time, but is clearly not a fundamental physical law; it can only be retrieved as an emergent, statistical property of macroscopic systems. As the largest conceivable macroscopic system, the universe itself becomes the natural context for understanding temporal phenomena. Our conjecture adopts the philosophical stance that the passage of time is not a fundamental feature of physics, but rather a cosmological effect, emerging from and intrinsically connected to the local and global structure of spacetime.
Our view departs from that of that of Castagnino and Lombardi [16] in that we insist that the arrow of time is defined by local thermodynamic physics, but that this can only be extended globally to the whole spacetime if the geometrical and topological conditions listed above are satisfied. Time is thus best understood as a cosmological phenomenon, having to do with the large scale mass and energy distribution in the universe, which gives rise to both the local thermodynamic time asymmetry (characterized by the Weyl curvature tensor), and the correct global structure that can support the asymmetry everywhere on the manifold.
While classically this is enough to give a reasonable instantiation of the A-theory ontology of time, quantum mechanics complicates matters. The fundamental equation of quantum cosmology, the Wheeler–DeWitt equation, is notoriously timeless, and reintroducing time in the description of a quantum system is an infamously difficult problem, more so than its classical counterpart. Many approaches to solving the “problem of time” have been tried in the past, each with their strengths and weaknesses. The one that is most akin to the perspective of this paper is the Tempus Post Quantum paradigm of [66], in which time is not present at the fundamental level, but can nevertheless emerge in the quantum regime. Specifically, we presented the semiclassical approach to the problem, in which time is set by the expansion of the universe and is thus tied to geometry. While this approach seemingly works, it does so only under the assumption that the universe is already large and classical, and thus cannot be taken as a full resolution of the problem. A full quantum gravity solution would explain how and why the precise geometrical and topological structure compatible with the existence of time emerges in the first place. This has yet to be explained.

### v3

@@ 1 Introduction
  • Local obstructions: The universe does not admit a well defined notion of cosmic time, or, if it does, it started in a high-entropy state, i.e. either Weyl’s postulate [24] or the past hypothesis [25] (or both) are violated. Geometrical obstructions of this type prevent even a local arrow of time from being defined.
  • Global obstructions: The universe is not time-orientable [26] or admits closed timelike curves (CTC) [27] that violate causality and along which entropy cannot globally increase, i.e. the universe is not chronological [28]. Causal or topological obstructions of this type prevent a global arrow of time from being defined everywhere on the spacetime manifold.
@@ 3 The Past Hypothesis
The situation is slightly different for accelerating spacetimes with $\Lambda>0$ [41, 42, 43]. In this case the SEC needs to be replaced by the weaker condition
$R_{\mu\nu}u^{\mu}u^{\nu}\geq 3H^{2}g_{\mu\nu}u^{\mu}u^{\nu},$ (3.7)
where $H^{2}=\Lambda/3$. As we saw, the metric admits a preferred slicing with timelike four-velocity field $u^{\mu}=(1,0,0,0)$, therefore (3.7) is simply
$R_{00}\geq-3H^{2}.$ (3.8)
Then, (3.4) gives the bound
$\dot{\theta}\leq-\frac{\theta^{2}}{3}+3H^{2}.$ (3.9)
The right-hand side is now negative only if $|\theta|$ is large enough. In fact, focusing happens only when
$\theta|_{\Sigma}<-3H,$ (3.10)
everywhere on a compact Cauchy hypersurface $\Sigma$.
What further constraint does this place on the metric (2.3)? The conditions (3.8) and (3.10) effectively select a subset of all possible cosmic-time metrics (2.3). To see this, let us consider the simple case of a homogeneous and isotropic universe. The line-element is given then by the FRW metric:
@@
where $a(t)>0$ is the scale factor and k=1,0,1k=-1,0,1 the curvature parameter. Now, one can show that for the metric (3.11), $R_{00}=-3\ddot{a}/a$, therefore $R_{00}\geq-3H^{2}$ whenever $\ddot{a}\leq H^{2}a$ for all values of $t$. Effectively, this selects universes in which the acceleration is never larger than the de Sitter value.
Exotic matter or quantum effects that violate the assumptions above can give rise to cosmological models that have no beginning. For example, the simple bouncing model described by
$a(t)=a_{b}\left[1+\left(\frac{t}{t_{b}}\right)^{2}\right]^{q},\qquad q>0,$ (3.12)
can be shown to arise due to quantum cosmological effects described by the Wheeler-DeWitt equation [44] and has no cosmological singularity. In fact, in this model $\ddot{a}\leq 0$ only when $|t/t_{b}|\geq 1/\sqrt{1-2q}$ with $0<q<1/2$. This means that around the bounce, i.e. for $|t/t_{b}|\leq 1/\sqrt{1-2q}$, the universe reverses course, starts expanding, and the singularity is avoided. In this region $\ddot{a}/a=2q/t_{b}^{2}\gg\Lambda$, so condition (3.8) is violated.
@@
$g_{ij}(t,\vec{x})=a^{2}(t)\left[\gamma_{ij}^{(0)}(\vec{x})+t^{n}\gamma_{ij}^{(1)}(\vec{x})+...\right],\qquad n>0,$ (3.14)
@@
To summarize, if conditions (2.12), (3.8), and (3.10) are satisfied, and $g_{ij}(t,\vec{x})\sim a^{2}(t)\gamma_{ij}^{(0)}(\vec{x})$ for $t\rightarrow 0$, the universe admits a notion of cosmic time with a well-defined direction (that orthogonal to the hypersurfaces of constant cosmic time) and orientation (proceeding from the low-entropy initial singularity). While this ensures the existence of an arrow of time on a local patch, there could still be topological effects that spoil chronology globally, as we will see in the next two sections.
@@ 4 Time-orientability
$(t,\vec{x})\sim\left((-1)^{n}t,\vec{x}+n\vec{a}\right)\qquad\forall n\in\mathbb{Z}.$ (4.3)
@@ 5 Chronology
$\gamma:\qquad r=r_{0}=\text{const},\qquad z=\text{const},\qquad\tau=\tau_{0}=\text{const},\qquad\phi\in[0,2\pi].$ (5.5)
@@ 6 The Role of Entropy
The first describes an expanding universe, while the second describes a contracting universe. Both are perfectly legitimate solutions of the dynamical equations, but only one of them can be physically realized [^20] Note again that it is the singularity at $t=0$ that prevents the two solutions from being glued together to restore symmetry, reinforcing the idea that a singularity is needed to define an arrow of time, see §3.. Whichever one chooses, time-reversal invariance will be broken.
@@
In §3 we argued that the entropy requirement can be reformulated as a geometric condition: the Weyl curvature hypothesis. This does not contradict the idea that the arrow of time arises from the second law. Rather, it reflects the fact that the entropy of the gravitational field is inherently geometric, since gravity itself is described by a geometric theory. Time is fundamentally tied to the large-scale geometry and topology of the universe, which must allow for a global asymmetry in the cosmic time direction. However, it is not just any geometric asymmetry of spacetime that will yield an arrow of time, but only the particular asymmetry encoded in one specific geometric object (the Weyl curvature tensor) which has a direct link to the entropy of the gravitational field.
There is another reason to believe that entropy must enter somehow into the definition of the arrow of time and this has to do with memory [11, 58]: the formation, stability, and accessibility of memories, whether in physical records, biological systems, or cognitive processes, require irreversible dynamics driven by an entropy gradient. The possibility of encoding and retrieving information about past states depends on the dissipation of free energy into heat, so that in a perfectly reversible world, no memory could persist. This ties the thermodynamic arrow directly to the psychological arrow of time: our perception of temporal flow, grounded in the accumulation of records of the past but not the future, emerges from the same entropic asymmetry that shapes the universe at large. Without increasing entropy, neither physical traces nor subjective experience of time’s direction could exist [^22] Castagnino and Lombardi argue instead that, once a global temporal asymmetry is found, the past-future distinction is conventional and can be expressed locally as a four-dimensional energy flow, without invoking entropy. We agree this is a coherent view, but maintain that entropy is essential to connect global asymmetry with irreversibility, memory, and the experienced passage of time..
@@ 7 The Problem of Time
$\mathcal{C}=0,\qquad\mathcal{C}^{i}=0.$ (7.13)
@@
$K_{ij}=\frac{1}{2N}\dot{h}_{ij}=\frac{a\dot{a}}{N}\gamma_{ij},\qquad K=\frac{3\dot{a}}{Na}.$ (7.19)
@@
$\pi^{ij}=\frac{-\sqrt{h}\dot{a}}{N\kappa a}h^{ij},\qquad\pi=-\frac{3\sqrt{h}\dot{a}}{N\kappa a}.$ (7.20)
@@ 8 Conclusions
The arrow of time is instead an emergent phenomenon set by the second law of thermodynamics. However, defining a consistent arrow of time on a generic spacetime manifold is only possible under very special circumstances, namely whenever the universe obeys certain geometrical and topological conditions:
@@
According to this view, what we understand as time (change, flux, becoming, etc.) is not tied in any way to fundamental physics, but is instead a product of the distribution of matter and energy in the whole universe, or, equivalently, since matter and energy determines geometry, of its shape. A simple example is given by irreversibility, which is one of the basic features of our experience of time, but is clearly not a fundamental physical law; it can only be retrieved as an emergent, statistical property of macroscopic systems. As the largest conceivable macroscopic system, the universe itself becomes the natural context for understanding temporal phenomena. Our conjecture adopts the philosophical stance that the passage of time is not a feature of microphysics, but rather a cosmological effect, emerging from and intrinsically connected to the local and global structure of spacetime.
While classically this is enough to give a reasonable instantiation of the A-theory ontology of time, quantum mechanics complicates matters. The fundamental equation of quantum cosmology, the Wheeler–DeWitt equation, is notoriously timeless, and reintroducing time in the description of a quantum system is an infamously difficult problem, more so than its classical counterpart. Many approaches to solving the “problem of time” have been tried in the past, each with their strengths and weaknesses. The one that is most akin to the perspective of this paper is the Tempus Post Quantum paradigm of [68], in which time is not present at the fundamental level, but can nevertheless emerge in the quantum regime. Specifically, we presented the semiclassical approach to the problem, in which time is set by the expansion of the universe and is thus tied to geometry. While this approach seemingly works, it does so only under the assumption that the universe is already large and classical, and thus cannot be taken as a full resolution of the problem. A full quantum gravity solution would explain how and why the precise geometrical and topological structure compatible with the existence of time emerges in the first place.